OpenStudy (anonymous):

Math help please and thank you :) Variable y varies directly with x^2, and y = 96 when x = 4. What is the value of y when x = 2? y=

OpenStudy (anonymous):

hello?

OpenStudy (displayerror):

Variable \(y\) varies directly with \(x^2\), meaning that as the values of\(x\) increases, the value of \(y\) also increases. Consider the example that they've provided you with. The function is given by \(y = Ax^2\), where \(A\) is some constant (it doesn't change). What is the value of \(A\)? That's what we have to find out first, and that is why we are given the example. \[96 = A (4)^2\] Solve for \(A\) here. Then plug in the new information and solve for \(y\), only this time our \(x\) value is 2 (we're solving the question that is being asked) \[y = A(2)^2\]

OpenStudy (arabpride):

Don't forget to close this post if you're done. Thanks! :3